A circular dashboard of radius 2 m is at a distance of 20 m from you. you throw a dart at it and it hits the dart at some point Q in the circle . what is the probability that Q is closer to the centerof the circle than the periphery
1
Expert's answer
2011-02-18T08:01:44-0500
Probability= (area of inner circle)/(area of outercircle) i.e (r/2)2/r2 =1/4=0.25.